Solve the following:


1) 2x+y = 6y


2) 3x = 3(2y+1)

Grade:11

1 Answers

vikas askiitian expert
13 years ago

1)   2x+y  = 6y

taking log on both sides 

(x+y)log2 = ylog6 

log6 = log3*2 = log3+log2

(x+y)log2 = y(log2+log3)

 xlog2 = ylog3

  2x = 3y

solution is only one , (x,y) = (0,0)